OpenStudy (anonymous):

Solve 8^(x+2)=32^(2x+4)

OpenStudy (anonymous):

2^3(x+2)=2^5(2x+4)........2"s cancels out...we have 3(x+2)=5(2x+4).........opening the bracket with the co-efficicents 3x+6=10x+20.................collect like terms 3x-10x=20-6 -7x=14.............divide both side by -7 -7x/-7 = -14/7 x = -2

OpenStudy (anonymous):

Thanks